OpenStudy (anonymous):

Then answer the question below, using complete sentences. While there is no rule that says you must solve a two-step equation by adding or subtracting first, why is it recommended?

OpenStudy (precal):

because if you divide by a number first in a two step equation, you will create fractions example 3x+5=20

OpenStudy (precal):

if you divide first then you create x + (5/3)=(20/3)

OpenStudy (anonymous):

thank you alot and could you also answer this for me in a full sentance What does it mean to "balance" an equation?

OpenStudy (precal):

No, you can write your own full sentences. I will be happy to give you brief statements of key ideas. Balance is used in equations, we must maintain balance. that is why we add numbers to both sides or subtract numbers to both sides or multiple numbers to both sides or divide numbers to both sides. We are keeping the balance.....

OpenStudy (anonymous):

thank you you are very helpfull i really apreciate it

OpenStudy (precal):

anytime